OpenStudy (anonymous):

Solve the following equation...PLZ SHOW WORK (x - 3) /x = (x - 3) / (x - 6)

OpenStudy (anonymous):

(x - 3) /x = (x - 3) / (x - 6) (x-3)(x-6)=x(x-3) x^2-9x+18=x^2-3x -6x=-18 x=3